Sam has 6 rose bushes. He counted the flowers on each of them. There are 8, 2, 5, 4, 11 and 9. Find the Standard Deviation. Is X a “Usual” number of flowers? X is your last two digits of your CUNYFirst ID Number. (00 = 0, 01 = 1, …) here X= 84


1
Expert's answer
2021-10-11T15:54:52-0400

data= 8,2,5,4,11,9

mean(data) "=\frac{8+2+5+4+11+9}{6} =6.5"

standard deviation "= \sqrt{\frac{(8-6.5)^2 +(2-6.5)^2 +...+(11-6.5)^2 +(9-6.5)^2}{6}} = 3.095"

number of flowers considered unusual if it is not lies between

(mean(data)-3 "\times" standard deviation, mean(data)+3 "\times" standard deviation)

"=(6.5-3 \times 3.0957, 6.5+3 \times 3.0957) \\

=(-2.7871, 15.7871)"

x=84 >15.7871 hence number of flowers considered unusual